### v1

@@ 3.1 Vision for large-scale systems
Here $P_{s}$ is the power consumption of a switch. Large-scale neural simulations are communication bound, rather than compute bound [75]. Here, we assume that the transmission power dominates the computing power of the neurons [30], specifically for networks at scale.
@@
Current trends in the evolution of ethernet switches: 51.2-Tbps Co-Packaged Optics Ethernet Switch Platform is one of the recently launched products by multiple companies such as CISCO, BroadComm, etc [11]. CISCO’s G200 switch consists of 512 x 100GE Ethernet ports on one device. The Marvell Teralynx 51.2T Ethernet Switch consumes 1 watt per 100 gigabits-per-second of bandwidth [63]. Similarly, the Broadcomm’s Tomahawk 4 Switch consumed a total power of 450 W, producing 1.76 W per 100 gigabits-per-second of bandwidth [52]. Considering 1W/100Gbps, the power consumption of a 51.2 Tbps switch ($P_{s}$) will be 512 Watts. In recent technology, Co-packaged optics (CPO) has evolved as a solution to meet the growing demand for data movement. CPO with the ethernet switch ASICs presents distinct benefits in terms of bandwidth, size, weight, and power consumption [59, 53]. The power consumption of pluggable optical modules vary from 15 pJ/bit to 20 pJ/bit. However, the power consumption of CPO systems can be decreased to a range of 5 pJ/bit to 10 pJ/bit [59]. Recently, Broadcomm presented the performance data of TH5-Bailly CPO 51.2-Tbps Ethernet switch delivering 5W/800 Gbps interconnect power [66], which amounts to 6.25 pJ/bit. Assuming a CPO-based interconnect power density of 6.25 pJ/bit the total interconnect power consumption is computed as:
Pi=6.25pJ/bit51.2Tbps=320WPi=6.25pJ/bit*51.2\ Tbps=320W (47)
@@
The total power consumption of the network infrastructure ( switch + interconnect) from Eq. 46 is 832W.
@@
Neuron in a core: Assuming a spiking frequency of 10 Hz (usual metric reported for the human brain), average fan in and fanout of 1K, for n neurons with 4-bit quantized weights [76] (It can be inferred from the robustness analysis that 4 bits is sufficient to maintain competitive accuracy), we solve Eq. 34 to compute the total number of neurons in a core to be 71 B neurons with each port being shared by 277 M neurons. The system supports 71 trillion synapses. We focus on the simplistic case of K=1 and K = 10 networks for ease of implementation and $H_{in}=H_{out}$, such that the average incoming and output in traffic are matched.
@@
Memory required to store the routing table: The memory required to store the synaptic information scales with the number of synaptic connections the system supports. However, the current ethernet switches support a shared on-chip memory in the range of MB only [52]. To support 71 trillion synaptic connections, external memory access is imperative. The total memory required by this system is around 26.6 TB (using Eq. 35). These days, industrial switches and routers incorporate high bandwidth memory (HBM) as packet buffers to store the packets in case of excess traffic [77, 78]. High Bandwidth Memory (HBM) is a game-changing technology that allows high-bandwidth switches to be equipped with deep buffering capacity [78]. The state-of-the-art Ethernet switches are equipped with HBM, which serves as packet buffers in case of excess traffic [78, 77], though they are in the range of Gbytes [78]. Assuming the ethernet switch is equipped with a larger HBM capacity to store and communicate the synaptic information, we compute the power cost incurred in HBM memory access and signaling. The projected energy for DRAM access and interface energy was 4.5 pJ/bit in 16nm technology [79]. HBM has relatively good energy efficiency (about 3-5pJ/b) compared to other DRAMs [80, 79, 81, 82, 83]. For every event, destination port, destination pool address, and the delays of the h synaptic connections have to be accessed. The number of bits to be accessed is:
$m_{access}=hlog_{2}(w)+log_{2}(m/h)+log_{2}(n/m)\ bits$ (48)
@@
Queue capacity: To support large-scale scale simulation the total shared memory required to support the shaped queues at the output ports for $K=1$ implementation is 45.3 MB (using Eq. 39) if we consider the synaptic delay to be in the range of microseconds unit and $n_{t}$ to be of 32 bits. (The current latency of the ethernet switches is in the range of 100s of nanoseconds [52].) If we consider $T_{out}$ in the range of 50 microseconds, then the memory required for the shared queues is 160 MB (Eq. 36). The state-of-the-art Ethernet switches already support sufficient memory capacity for packet buffers [78, 52] in the range of 100MBs. For K=10 networks, 10x the memory capacity will be required, in which case, HBMs have to be used (current switches support GBytes of HBM storage).
Table 2: Trends in technological advancements for interconnect systems
@@
00footnotetext: SE - energy efficiency of switch, IE - energy efficiency of interconnect, HBME - energy efficiency of HBM, SP - switch throughput in Tbps
Table 3: Performance projection numbers for $\pi^{2}$ network implementation
@@
00footnotetext: $E$ - energy efficiency of the interconnect system assuming K=1.
@@ 4 Simulator implementation of $\pi^{2}$-NN using traffic shaping protocols
Figure 13: a) $\pi^{2}$ network simulation on OMNET++: This implementation exploits the existing functionalities of Ethernet routers for $\pi^{2}$ computation. The transmitter and receiver nodes are TSN devices that are compatible with the TSN protocols. Each TSN node receives the sensor inputs and generates events (frames) with the corresponding PCP tags (synaptic weights are stored in memory) and destination addresses at specific time instants. The input spikes from the Tx nodes are first routed to the TSN switch implementing the asynchronous time shaper (ATS) algorithm, and the membrane potential accumulation is implemented by another TSN switch, which follows credit-based traffic shaping (CBS). An instantiation of a 2x10x2 network is portrayed here for illustration purposes. Similarly, we run a 784x50x10 network using the simulator.
We present a simulation model of the proposed $\pi^{2}$ architecture on the OMNET++ platform which is a C++ framework that enables the construction of network simulators and has extensive libraries to support the TSN protocol [32]. OMNET++ already has built-in codes to demonstrate various traffic shaping mechanisms relevant to TSN networks. We implement a 784x50x10 $\pi^{2}$ network trained on the MNIST dataset [23] as a proof of concept.
@@
  • Node: Each node (Tx and Rx) consists of an input port, through which it receives event data from external sources (sensor) or other nodes in the system, a packet generation logic, and a local memory where it stores the synaptic information (destination addresses and synaptic delays). It outputs a packet with the appropriate destination address and a PCP tag (representing the delay value) attached to it. Usually, the Ethernet frames consist of a priority code point (PCP) tag, which is used to identify their traffic class [27]. The nodes are simulated as TSN (Time-sensitive networking) device components on OMNET++ (as portrayed in Fig. 13) [32].
@@
  • The input spikes from the Tx nodes are first routed to the ATS switch, implementing the asynchronous time shaper (ATS) algorithm. The main components of the ATS switch are portrayed in Fig. 14. Based on the PCP value of the frame, the classifier routes it through the appropriate gate. The ATS meter computes the TET value of the frame based on the PCP tag value. We modify the equation to compute the TET for each event, based on the PCP tag (synaptic delay) associated with each frame. So, if the event with a PCP tag value of $W_{i}$ enters the switch at time $T_{i}$, then the TET for the frame is computed as:
    $TET=T_{i}+W_{i}$ (49)
@@
    Usually, the PCP is 3-bit code and can support 8 different traffic classes [27]. Accordingly, each egress port of the switch consists of 8 queues. The events are dropped into the queues based on their PCP tag values, buffered until their TET is reached, and are transmitted to the ingress port of the TSN-CBS switch. Inherently, the events in the queues of the egress ports will be arranged in ascending order of their TET values.
@@
  • The main components of the CBS switch are portrayed in Fig. 15. Every egress port of the CBS switch is connected to a particular Rx node. Each of these egress ports is associated with two queues (to support differential computing). The traffic shaping algorithm decides when the gate for the queue opens for event transmission. The CBS switch implements a modified version of the credit-based shaping algorithm (Fig. 15). We modify the credit calculation algorithm to suit the $\pi^{2}_{K}$ neuron dynamics. The modified CBS algorithm computes the credits (membrane potential) at every local time instant with a slope defined by the number of events (maximum K packets are stored; the remaining are dropped due to congestion) stored in its corresponding queue. When the credit value reaches a threshold (gamma), the CBS gate of that queue opens, and a packet is directed from the queue to the corresponding Rx node. The remaining packets in the queue are dropped (non-linearity).
@@
Figure 14: The inner architecture of the ATS switch - Based on the PCP value (3-bit), the classifier routes the frame through the right gate. The ATS meter computes the TET value and allows frame transmission when the TET value is reached. The snapshot of this ATS mechanism at work between the input and hidden layer is portrayed here.
@@
Figure 15: The inner architecture of the CBS switch - The hidden layer is constrained to K=140. Thus, the shared queue (capacity) is configured to hold only 140 frames. The frames will inherently arrive in a time-sorted manner in the queues. In case of congestion, the other frames are dropped as displayed. Every hidden layer node receives 1568 (784x2) frames from the input layer. Out of which 140 are contained, one frame is transmitted, and the remaining are dropped due to congestion in the queue[0] and queue[1].
@@
We trained a 3-layer $\pi^{2}$ based network on the MNIST dataset, quantized the weights (delays) to 3 bits (to map to 3-bit PCP value), and performed inference on the OMNET++ simulator. This experiment gives an idea of an implementation strategy where the fundamental operations of the $\pi^{2}$ model, namely delaying, sorting (in TSN-ATS switch), event dropping, and thresholding (in TSN-CBS switch), can be mapped to existing traffic shaping protocols of Ethernet switches. This implementation on the simulator provides a simplistic view to demonstrate how Ethernet switches can be used for computation and communication simultaneously using traffic-shaping protocols.
@@ 5 Effect of weight distillation
A conventional MAC-based ResNet-9 network is first trained on the CIFAR-10 dataset, and its weights are then ported to the $\pi^{2}$-ResNet-9 architecture. The $\pi^{2}$ network is subsequently fine-tuned to achieve the baseline competitive accuracy. However, when weight distillation strategy is omitted the $\pi^{2}$- network experiences an accuracy drop of approximately 1-2%.
Figure 16: The impact of weight porting (distillation) (WP) on training dynamics and generalization performance of the ResNet-9 network trained on the CIFAR-10 dataset is illustrated. (A,B) Test accuracy (%) and training loss over training epochs with and without weight distillation from a trained conventional teacher network.

### v3

@@ 3.2 Vision for large-scale systems
Here, $P_{s}$ is the power consumption of a switch. Large-scale neural simulations are communication-bound, rather than compute-bound [spinnaker2]. Here, we assume that the transmission power dominates the computing power of the neurons [stanojevic2024], specifically for networks at scale.
@@
Current trends in the evolution of Ethernet switches: 51.2-Tbps Co-Packaged Optics Ethernet Switch Platform is one of the recently launched products by multiple companies such as CISCO, Broadcom, etc [cisco]. CISCO’s G200 switch consists of 512 x 100GE Ethernet ports on one device. The Marvell Teralynx 51.2T Ethernet Switch consumes 1 watt per 100 gigabits-per-second of bandwidth [marvel]. Similarly, the Broadcom’s Tomahawk 4 Switch consumed a total power of 450 W, producing 1.76 W per 100 gigabits-per-second of bandwidth [25.6]. Considering 1W/100Gbps, the power consumption of a 51.2 Tbps switch ($P_{s}$) will be 512 Watts. In recent technology, Co-packaged optics (CPO) has evolved as a solution to meet the growing demand for data movement. CPO with the Ethernet switch ASICs presents distinct benefits in terms of bandwidth, size, weight, and power consumption [cpoprogress, cpo]. The power consumption of pluggable optical modules varies from 15 pJ/bit to 20 pJ/bit. However, the power consumption of CPO systems can be decreased to a range of 5 pJ/bit to 10 pJ/bit [cpoprogress]. Recently, Broadcom presented the performance data of TH5-Bailly CPO 51.2-Tbps Ethernet switch delivering 5W/800 Gbps interconnect power [broadcommcpo], which amounts to 6.25 pJ/bit. Assuming a CPO-based interconnect power density of 6.25 pJ/bit, the total interconnect power consumption is computed as:
Pi=6.25pJ/bit51.2Tbps=320WPi=6.25pJ/bit*51.2\ Tbps=320W (S28)
@@
The total power consumption of the network infrastructure ( switch + interconnect) from Eq. S27 is 832W.
@@
Neuron in a core: Assuming a spiking frequency of 10 Hz (usual metric reported for the human brain), average fan in and fanout of 1K, for n neurons with 4-bit quantized weights [quantize] (It can be inferred from the robustness analysis that 4 bits is sufficient to maintain competitive accuracy), we solve Eq. S15 to compute the total number of neurons in a core to be 71 B neurons with each port being shared by 277 M neurons. The system supports 71 trillion synapses. We focus on the simplistic case of K=1 and K = 10 networks for ease of implementation and $H_{in}=H_{out}$, such that the average incoming and outgoing traffic are matched.
@@
Memory required to store the routing table: The memory required to store the synaptic information scales with the number of synaptic connections the system supports. However, the current Ethernet switches support a shared on-chip memory in the range of MB only [25.6]. To support 71 trillion synaptic connections, external memory access is imperative. The total memory required by this system is around 35.5 TB (using Eq. S16). These days, industrial switches and routers incorporate high bandwidth memory (HBM) as packet buffers to store the packets in case of excess traffic [broadcommHBM, ciscohbm]. High Bandwidth Memory (HBM) is a game-changing technology that allows high-bandwidth switches to be equipped with deep buffering capacity [ciscohbm]. The state-of-the-art Ethernet switches are equipped with HBM, which serves as packet buffers in case of excess traffic [ciscohbm, broadcommHBM], though they are in the range of Gbytes [ciscohbm]. Assuming the Ethernet switch is equipped with a larger HBM capacity to store and communicate the synaptic information, we compute the power cost incurred in HBM memory access and signaling. The projected energy for DRAM access and interface energy was 4.5 pJ/bit in 16nm technology [hbm_energy2]. HBM has relatively good energy efficiency (about 3-5pJ/b) compared to other DRAMs [hmb_energy3, hbm_energy2, hbm_energy4, hbm_energy5, hbm_energy6]. For every event, the destination port, destination pool address, and the delays of the h synaptic connections have to be accessed. The number of bits to be accessed is:
$m_{access}=hp+log_{2}(m/h)+log_{2}(n/m)\ bits$ (S29)
@@
Queue capacity: To support large-scale simulation, the total shared memory required to support the shaped queues at the output ports for $K=1$ implementation is 45.3 MB (using Eq. S20) if we consider the synaptic delay to be in the range of microseconds and $n_{t}$ to be of 32 bits. (The current latency of the Ethernet switches is in the range of 100s of nanoseconds [25.6].) If we consider $T_{out}$ in the range of 50 microseconds, then the memory required for the shared queues is 160 MB (Eq. S17). The state-of-the-art Ethernet switches already support sufficient memory capacity for packet buffers [ciscohbm, 25.6] in the range of 100MBs. For K=10 networks, 10x the memory capacity will be required, in which case, HBMs have to be used (current switches support GBytes of HBM storage).
Table S2: Trends in technological advancements for interconnect systems00footnotetext: SE - energy efficiency of switch, IE - energy efficiency of interconnect, HBME - energy efficiency of HBM, SP - switch throughput in Tbps
@@
Table S3: Performance projection numbers for $\pi^{2}$ network implementation00footnotetext: $E$ - energy efficiency of the interconnect system assuming K=1.
@@
3.3 Analyzing the effects of Ethernet header overheads and higher synaptic bit precisions
[…]
If we consider the header, it will have a direct impact on the neuron count and performance projections. Specifically, Eq. S15 gets modified to:
$B\geq 512F_{spk}m$ (S30)
Solving Eq.S30 with $F_{spk}=10Hz$ and 100 Gbps bandwidth leads to $m=19.5M$ neurons sharing a port (14x reduction). Accordingly, the total neuron count reduces to $5B$ neurons, and the total memory required to store the look-up tables reduces to 2.5 TB. Following a similar analysis as above, we report the energy, throughput and performance projections with the header constraint in Table S4.
Table S4: Performance projection numbers for $\pi^{2}$ network implementation with Ethernet header overhead00footnotetext: $E$ - energy efficiency of the interconnect system assuming K=1.
Year Synapses(T) Throughput (petaOp/s) $P_{total}$ (kW) EE(pJ/Op)
2020 2.5 0.025 0.675 29
2022 5 0.05 0.83 17.8
2034 400 4 8.2 2.44
[…]
@@ 4 Simulator implementation of $\pi^{2}$-NN using traffic shaping protocols
Figure S6: a) $\pi^{2}$ network simulation on OMNET++: This implementation exploits the existing functionalities of Ethernet routers for $\pi^{2}$ computation. The transmitter and receiver nodes are TSN devices that are compatible with the TSN protocols. Each TSN node receives the sensor inputs and generates events (frames) with the corresponding PCP tags (synaptic weights are stored in memory) and destination addresses at specific time instants. The input spikes from the Tx nodes are first routed to the TSN switch implementing the asynchronous time shaper (ATS) algorithm, and the membrane potential accumulation is implemented by another TSN switch, which follows credit-based traffic shaping (CBS). An instantiation of a 2x10x2 network is portrayed here for illustration purposes. Similarly, we run a 784x50x10 network using the simulator.
We present a simulation model of the proposed $\pi^{2}$ architecture on the OMNET++ platform, which is a C++ framework that enables the construction of network simulators and has extensive libraries to support the TSN protocol [omnet_tsn]. OMNET++ already has built-in code to demonstrate various traffic shaping mechanisms relevant to TSN networks. We implement a 784x50x10 $\pi^{2}$ network trained on the MNIST dataset [temp] as a proof of concept.
@@
  • Node: Each node (Tx and Rx) consists of an input port, through which it receives event data from external sources (sensor) or other nodes in the system, a packet generation logic, and a local memory where it stores the synaptic information (destination addresses and synaptic delays). It outputs a packet with the appropriate destination address and a PCP tag (representing the delay value) attached to it. Usually, the Ethernet frames consist of a priority code point (PCP) tag, which is used to identify their traffic class [nesting]. The nodes are simulated as TSN (Time-sensitive networking) device components on OMNET++ (as portrayed in Fig. S6) [omnet_tsn].
@@
  • The input spikes from the Tx nodes are first routed to the ATS switch, implementing the asynchronous time shaper (ATS) algorithm. The main components of the ATS switch are portrayed in Fig. S7. Based on the PCP value of the frame, the classifier routes it through the appropriate gate. The ATS meter computes the TET value of the frame based on the PCP tag value. We modify the equation to compute the TET for each event, based on the PCP tag (synaptic delay) associated with each frame. So, if the event with a PCP tag value of $W_{i}$ enters the switch at time $T_{i}$, then the TET for the frame is computed as:
    $TET=T_{i}+W_{i}$ (S31)
@@
    Usually, the PCP is a 3-bit code and can support 8 different traffic classes [nesting]. Accordingly, each egress port of the switch consists of 8 queues. The events are dropped into the queues based on their PCP tag values, buffered until their TET is reached, and then transmitted to the ingress port of the TSN-CBS switch. Inherently, the events in the queues of the egress ports will be arranged in ascending order of their TET values.
@@
  • The main components of the CBS switch are portrayed in Fig. S8. Every egress port of the CBS switch is connected to a particular Rx node. Each of these egress ports is associated with two queues (to support differential computing). The traffic shaping algorithm decides when the gate for the queue opens for event transmission. The CBS switch implements a modified version of the credit-based shaping algorithm (Fig. S8). We modify the credit calculation algorithm to suit the $\pi^{2}_{K}$ neuron dynamics. The modified CBS algorithm computes the credits (membrane potential) at every local time instant with a slope defined by the number of events (maximum K packets are stored; the remaining are dropped due to congestion) stored in its corresponding queue. When the credit value reaches a threshold (gamma), the CBS gate of that queue opens, and a packet is directed from the queue to the corresponding Rx node. The remaining packets in the queue are dropped (non-linearity).
@@
Figure S7: The inner architecture of the ATS switch - Based on the PCP value (3-bit), the classifier routes the frame through the right gate. The ATS meter computes the TET value and allows frame transmission when the TET value is reached. The snapshot of this ATS mechanism at work between the input and hidden layer is portrayed here.
@@
Figure S8: The inner architecture of the CBS switch - The hidden layer is constrained to K=140. Thus, the shared queue (capacity) is configured to hold only 140 frames. The frames will inherently arrive in a time-sorted manner in the queues. In case of congestion, the other frames are dropped as displayed. Every hidden layer node receives 1568 (784x2) frames from the input layer. Out of which 140 are contained, one frame is transmitted, and the remaining are dropped due to congestion in the queue[0] and queue[1].
Figure S9: A) A 784x50x10 network is simulated using the proposed traffic shaping protocols and Ethernet switches on the OMNETPP software. B,C,D) The spatio-temporal patterns for input of different classes are extracted from the simulator and portrayed in the figures. The index of the node that generates a packet/frame first (earliest) denotes the class the input image belongs to. The output spike times exactly match the software simulations of the quantized $\pi^{2}$ network.
@@
We trained a 3-layer $\pi^{2}$ based network on the MNIST dataset, quantized the weights (delays) to 3 bits (to map to 3-bit PCP value), and performed inference on the OMNET++ simulator. This experiment gives an idea of an implementation strategy where the fundamental operations of the $\pi^{2}$ model, namely delaying, sorting (in TSN-ATS switch), event dropping, and thresholding (in TSN-CBS switch), can be mapped to existing traffic shaping protocols of Ethernet switches. The spiking times of the 10 output nodes for different images are extracted and used for classification as shown in Fig. S9. This implementation on the simulator provides a simplistic view to demonstrate how Ethernet switches can be used for computation and communication simultaneously using traffic-shaping protocols.
@@ 5 Effect of weight distillation
A conventional MAC-based ResNet-9 network is first trained on the CIFAR-10 dataset, and its weights are then ported to the $\pi^{2}$-ResNet-9 architecture. The $\pi^{2}$ network is subsequently fine-tuned to achieve the baseline competitive accuracy. However, when the weight distillation strategy is omitted, the $\pi^{2}$- network experiences an accuracy drop of approximately 1-2%.
Figure S10: The impact of weight porting (distillation) (WP) on training dynamics and generalization performance of the ResNet-9 network trained on the CIFAR-10 dataset is illustrated. (A,B) Test accuracy (%) and training loss over training epochs with and without weight distillation from a trained conventional teacher network.